find the dimension of a rectangular lot whose area is 3600sq. m. and whose length exceeds its width by 50m.?

Answer by mananth(16949)

width be x
length = x+50
Area = 3600
Area = L*W
x(x+50)=3600
x^2+50x=3600
x^2+50x-3600=0
x^2+90x-40x-3600=0
x(x+50)-40(x+50)=0
(x-40)(x+50)=0
Take the positive value for dimension
width = 40 m
length = 40+50 = 90 m
..
Check
Area = 90*40=3600m^2
m.ananth@hotmail.ca